OpenStudy (anonymous):

What is the solution to y = 0.7x for the replacement set x = {2}?

OpenStudy (anonymous):

@Reptile_Girl777

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

y = 0.7(2) what is 0.7 times 2?

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

Correct

OpenStudy (jackthegreatest):

yep

OpenStudy (anonymous):

y = 1.4 when x = 2

OpenStudy (anonymous):

the answer hole thing says ( , )

OpenStudy (anonymous):

Yes, (2, 1.4)

OpenStudy (jackthegreatest):

MDoodler is right

OpenStudy (anonymous):

so its (2,1.4)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the solution to y = 2x – 1 for the replacement set x = {1.4}?

OpenStudy (anonymous):

is is 1.8 the answer

OpenStudy (anonymous):

y = 2(1.4) – 1 y = 2.8 - 1 y = 1.8 Yes correct

OpenStudy (anonymous):

so (1.4,1.8)

OpenStudy (anonymous):

(1.4, 1.8)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thx so much